OpenStudy (anonymous):

Simplify the expression. 4(x+1)+2(x-1)

OpenStudy (skullpatrol):

Any ideas @paischil1 ? Good ideas make good answers?

OpenStudy (anonymous):

nope. having a really hard time with this one..

OpenStudy (skullpatrol):

$$\huge \text{Recall the distributive property:}$$ $$\huge a(b+c)=ab+ac$$ $$\huge a(b-c)=ab-ac$$

OpenStudy (skullpatrol):

$$\huge 4(x+1)+2(x-1)$$

OpenStudy (anonymous):

6x+2

OpenStudy (skullpatrol):

Yes. Nice work :)